OpenStudy (anonymous):

Solve the quadratic equation y2 - 5y = -4 by factoring.

OpenStudy (anonymous):

first you'll need to put it in standard form: \(\large y^2-5y+4=0 \) now, think what two numbers, r1 and r2 such that: r1 + r2 = -5 AND r1 * r2 = 4 can you think of any ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

you can't think of any two numbers that multiply to 4?